OpenStudy (anonymous):

the perimeter of a triangle is twice the sum of its length and width the perimeter is 34 meters and its length is 2 meters more than twice the width. what is the length?

OpenStudy (anonymous):

p*2=2(l+2) p=34 l=2w+2 solve that system of equations

OpenStudy (anonymous):

p*2=2(l+2), p=34, l=2w+2 l=32

OpenStudy (anonymous):

@timo86m Thankyou!